OpenStudy (kayders1997):

Find an equation ion to the normal line to the graph of y=(3x^2+2x)^1/2 at (2,4) is

OpenStudy (kayders1997):

@zepdrix

OpenStudy (kayders1997):

Ugh?

OpenStudy (kayders1997):

Okay?

OpenStudy (kayders1997):

@mathmale

zepdrix (zepdrix):

Do you remember how to find `tangent line`?

OpenStudy (kayders1997):

Take the derivative right?

OpenStudy (kayders1997):

Maybe o_O

zepdrix (zepdrix):

at the point (2,4), f'(2) gives us the slope of the tangent line. The slope of the normal line will be the `negative reciprocal of the tangent slope`.

zepdrix (zepdrix):

\[\large\rm m=-\frac{1}{f'(2)}\]

zepdrix (zepdrix):

So yes, start by finding a derivative.

OpenStudy (kayders1997):

Okay

OpenStudy (kayders1997):

Do you want it in square root form?

OpenStudy (kayders1997):

Well 1/2(6x+2)/square root 6x^2+2x

OpenStudy (kayders1997):

Omg :/ of course

OpenStudy (kayders1997):

She's here, her aunt usually takes 10 hours to shop

zepdrix (zepdrix):

\[\large\rm f'(x)=\frac{1}{2}(3x^2+2x)^{-1/2}(6x+2)\]Ok good. Evaluate at x=2.

zepdrix (zepdrix):

oh lol :P ok peace out

OpenStudy (kayders1997):

We can just keep this and than if we're on at the same time one of us can tag eachother

zepdrix (zepdrix):

ya

OpenStudy (kayders1997):

@zepdrix when you get on :D

zepdrix (zepdrix):

\[\large\rm f'(x)=\frac{1}{2}(3x^2+2x)^{-1/2}(6x+2)\]Evaluate at x=2,\[\large\rm f'(2)=\frac{1}{2}(3(2)^2+2(2))^{-1/2}(6(2)+2)\]Simplify. This is the tangent slope. Flip it, negative it, this is the normal slope.

OpenStudy (kayders1997):

@zepdrix

OpenStudy (kayders1997):

7/4?

OpenStudy (kayders1997):

So -4/7? For normal line

zepdrix (zepdrix):

Good good good, that is our `normal slope`.

zepdrix (zepdrix):

\[\large\rm m=-\frac{4}{7}\]

zepdrix (zepdrix):

When you've done these problems in the past, do you usually set them up in `point-slope form` or `slope-intercept form`?

OpenStudy (kayders1997):

Point slope

OpenStudy (kayders1997):

Usually

zepdrix (zepdrix):

Ok good, then we're done.

zepdrix (zepdrix):

\(\large\rm y-y_o=m(x-x_o)\) Just plug in the stuff.

zepdrix (zepdrix):

\(\large\rm (x_o,y_o)=(2,4)\)\[\large\rm m=-\frac47\]

OpenStudy (kayders1997):

Y-4=-4/7(x-2)

zepdrix (zepdrix):

yay team

OpenStudy (kayders1997):

:)

OpenStudy (kayders1997):

So?